OpenStudy (anonymous):

If an equation is in standard form how do you change it to y-intercept form If an equation is in standard form how do you change it to y-intercept form @Mathematics

OpenStudy (chriss):

solve for y

OpenStudy (chriss):

Ax+By=c and you want to get y=mx+b Ax+By=c By=-Ax+c y=(-Ax+c)/B m=-A/B and b = c/B